OpenStudy (anonymous):

can we simplify 2t^2/2t

OpenStudy (amistre64):

2 t t ---- 2 t so maybe

OpenStudy (mani_jha):

Yeah, why not. See that there it t common in both the numerator and denominator. Just remove a t from each side, and also cancel the 2. 2t^2/2t=t^2/t(cancelling 2)=(t*t)/t=t

OpenStudy (amistre64):

its not so much of a cancel; thats just the verbial equivalent of the actual mathing that goes on anything divided by itself equals 1 and 1 times anything else does not change the value of the expression so simply ignoring the 1s as superfluous is fine

OpenStudy (anonymous):

can u all be certain of the results?

OpenStudy (amistre64):

at this point no, since we are simply assuming that your notation corresponds to a normative mathematical design and not some misconstrued concoction of your own making...

OpenStudy (anonymous):

ok express w in terms of d and s s=\[s=\sqrt{3d(w-d)/8}\]

OpenStudy (amistre64):

square, times 8, divide off the 3d and add a d

OpenStudy (anonymous):

can u do it using equations?

OpenStudy (amistre64):

most likely, but then so could you if you follow those instructions :)

OpenStudy (anonymous):

i knw i got the answer, in a way d^2 / d, so can v like cncel it of..

OpenStudy (amistre64):

d/d = 1 so that leaves you with a: d

OpenStudy (anonymous):

can u solve t nd give me the answer